Solve the following equations: 
 2y + (5/2) = (37/2)............

Grade:12

2 Answers

Harshit Singh
3 years ago
Dear Student

By transposing (5/2) from LHS to RHS it becomes -5/2
Then,
= 2y = (37/2)–(5/2)
= 2y = (37-5)/2
= 2y = 32/2
Now,
Divide both side by 2,
= 2y/2 = (32/2)/2
= y = (32/2) × (1/2)
= y = 32/4
= y =8

Thanks
Shraddha Singh
2 years ago
Dear Student,
 
We can solve it as follows :
2y + 5/2 = 37/2
2y = 37/2 - 5/2
2y = (37-5)/2
2y = 32/2
2y = 16
y = 8
 
Hope it helps!
Thank you ✌️